# Are Vision xLSTM Embedded $U$-Nets More Reliable in Medical 3D Image Segmentation?

### v2

@@ Abstract
The development of efficient segmentation strategies for medical images has evolved from its initial dependence on Convolutional Neural Networks (CNNs) to the current investigation of hybrid models that combine CNNs with Vision Transformers. There is an increasing focus on creating architectures that are both high-performance and computationally efficient, able to be deployed on remote systems with limited resources. Although transformers can capture global dependencies in the input space, they face challenges from the corresponding high computational and storage expenses involved. This paper investigates the integration of CNNs with Vision Extended Long Short-Term Memory (Vision-xLSTM)s by introducing the novel U-VixLSTM.
The Vision-xLSTM blocks capture temporal and global relationships within the patches, as extracted from the CNN feature maps. The convolutional feature reconstruction path upsamples the output volume from the Vision-xLSTM blocks, to produce the segmentation output. Our primary objective is to propose that Vision-xLSTM forms an appropriate backbone for medical image segmentation, offering excellent performance with reduced computational costs. The U-VixLSTM exhibits superior performance, compared to the state-of-the-art networks in the publicly available Synapse, ISIC and ACDC datasets. Code provided: https://github.com/duttapallabi2907/U-VixLSTM
@@ 1 Introduction
Artificial intelligence-driven medical image segmentation is one of the key cornerstones of modern healthcare, providing essential information for diagnosing, planning treatment, and monitoring patients. This task includes both volumetric and 2D imaging modalities. There are several challenges in accurately defining target regions from medical image data. These challenges stem from the complex characteristics of medical imaging data and variations in anatomical structures. Some common issues include ambiguous boundaries of target organs and pathologies, low contrast, overlapping structures, and inter- and intra-patient variability. Volumetric medical image data involves examining the spatial relationship between regions of interest (ROI) in numerous slices of volumetric data [1]. Consequently, it requires a substantial amount of computational resources for analysis. These challenges require the development of segmentation algorithms with a high degree of precision and accuracy at the output.
@@
Deep learning is widely used in automated medical image segmentation, mainly due to its ability to extract underlying characteristics from input images with minimal human involvement and generalization to unseen samples [2]. The $U-$Net architecture [3], consisting of symmetric encoder-decoder structures with CNN, played a crucial role in revolutionizing the extensive use of deep learning in medical image segmentation. Hierarchical modeling of complex high-level patterns from low-level features and the integration of fine-grained spatial information with coarser abstract features are the factors that are attributed to the success of the $U-$Net. Various well-known architectures have been developed to expand on their accomplishments. These include $U-$Net++ [4], $V-$Net [5], Attention $U-$Net [6] and $U-$Net 3+ [7]. Each of these introduces distinct modifications to improve performance and address specific difficulties inherent in medical image segmentation.
@@
Combinations of CNNs and Transformers have been explored to combine their generic benefits to create a robust representation of target structures. Transformers overcome a significant drawback of CNNs through their self-attention mechanism; thereby, enabling them to effectively capture global relationships. In contrast, CNNs focus mainly on local patterns because of their limited field of view. TransUNet [8] pioneered the study of the impact of such a hybrid structure on medical image segmentation. It modeled the global structural dependencies in the intermediate feature volumes, produced from CNN layers, using the self-attention mechanism of Transformers [9]. This framework was also adopted into multiple segmentation tasks in the medical domain with the necessary modifications [10], [11] and [12]. Some other well-known hybrid CNN-Transformer architectures in this domain include UNETR [13], Swin UNETR [14], TransAttUNET [15], UNETR++ [16] and DS-TransUNet [17].
@@
The self-attention mechanism, which is the driving force behind Transformers, suffers from the curse of quadratic complexity. This leads to a huge computational burden and memory requirement. Thereby, the applicability of Transformers gets limited in resource-constrained environments. Extended Long-Short Term Memory (xLSTM) [18] has recently emerged as a strong contender to Transformers in sequence modeling. It addresses the key limitations of Transformers with its inherent linear computational complexity [$\mathcal{O}(n)$] and constant memory complexity [$\mathcal{O}(1)$] w.r.t. the length of input sequences. Vision-xLSTM [19] pioneers the application of xLSTMs to computer vision tasks.
The efficiency of deep algorithms entails goodness of performance, both in terms of resource consumption and accuracy. This ensures the scalability of AI-driven algorithms in real-world problems. Clinical scenarios, with limited computing resources, require models that can accurately delineate target structures from medical data while constraining computational power and memory requirements. This is essential to guarantee sustainable and cost-effective healthcare solutions.
@@
This need for computationally efficient and highly accurate segmentation algorithms motivated us to develop the new U-VixLSTM model for medical image segmentation. Our approach integrates CNNs with Vision-xLSTMs for the first time (to the best of our knowledge) for segmentation. The architecture is based on the popular $U-$ shaped framework reported in the literature [20], [21], [22], [23]. Our model is developed in two different versions: one for handling volumetric inputs and the other for handling 2D medical images. Our research contribution is summarized below.
@@
  • CNNs in the feature extraction path initially capture fine-grained textural information and local patterns corresponding to the target anatomical structures from the input image. The Vision-xLSTM blocks encode the global context within the intermediate output volumes, as obtained from the CNN layers.
@@
  • The feature reconstruction path upsamples the output from Vision-xLSTM blocks to produce the final segmentation output. Skip connections concatenate the feature volume from each level of the feature extraction path to the feature volume of the corresponding level of the feature reconstruction path, to build a robust representation of the target structures.
@@
The remaining sections of the paper are structured in the following manner. Section 2 discusses the existing literature. Section 3 presents the detailed description of the proposed architecture U-VixLSTM. The experimental results are illustrated in Section 4, along with a comparative study with state-of-the-art networks on publicly accessible datasets to validate the efficacy of U-VixLSTM. Finally, the article is concluded in Section 5.
Figure 1: Architectural framework of U-VixLSTM depicting the input processed by stacked layers of CNNs and Vision-xLSTM blocks. The intermediate feature representation from the feature extraction path is upsampled through the feature reconstruction path to obtain the final segmentation output.
2 Background Literature
TransUNet [8] introduced the pioneering approach to combine CNNs and Transformers to leverage their distinct benefits. The TransUNet framework was then improved [10] by incorporating an auxiliary decoder arm. A two-stage deep learning framework was used [11] for multi-organ segmentation from pelvic CT images. The initial phase involved the localization of organs utilizing a hybrid encoder framework, which comprised a convolutional stage followed by transformer modules. This output was subjected to several convolution operations.
Recent advances in attention mechanisms and transformer-based architectures have significantly improved the performance of medical image segmentation tasks. An efficient attention mechanism [27] was introduced to combine linear self-attention with frequency attention, using a Laplacian pyramid to selectively emphasize frequency contributions. This approach employs a fusion strategy to balance the shape and texture characteristics while improving overall segmentation accuracy. UCTransNet [28] integrates a multi-scale channel cross-fusion transformer (CCT) and channel cross-attention (CCA) to enhance medical image segmentation. The CCT module fuses multiscale features through cross-attention mechanisms to capture long-range dependencies, while the CCA aids in refining Transformer and decoder feature fusion. This methodology has demonstrated a significant improvement in segmentation performance on multiple benchmarks. The TransAttU-Net [15], a Transformer-based attention-guided $U$-Net, integrates multilevel guided attention and multiscale skip connections for biomedical image segmentation. The Transformer Self-Attention and Global Spatial Attention mechanisms capture long-range dependencies and spatial relationships, while multiscale skip connections improve feature aggregation across semantic scales. This results in improved segmentation accuracy and consistency. An efficient Vision Transformer [29] was designed for medical image segmentation, featuring a Mixed Transformer module that includes Local-Global Self-Attention (LGG-SA) and External Attention (EA). While the LGG-SA captures local and global dependencies, using a focalized computation strategy, the EA exploits inter-sample correlations. This design reduced computational complexity, while improving segmentation accuracy, compared to traditional transformers. Recent advancements involve synergizing Mamba blocks with the shifting window approach of Swin UNETR, aiming to harness the benefits of State Space Models and Transformers [30].
Although the advanced methodologies listed can yield significant performance improvements, it is crucial to recognize their accompanying complexities. Complex operations often result in higher memory and computational demands. The concurrent matrix computations across various attention heads, typical of these methodologies, require considerable processing power. This may lead to delayed convergence during training, requiring extended training durations. They potentially impede scalability, particularly when processing high-resolution volumetric medical images. Incorporating auxiliary modules, although improving performance, inevitably increases the size and complexity of the model. This can present challenges in resource-constrained environments while increasing the risk of overfitting.
In this context, our study explores the hybridization of CNNs and Vision-xLSTMs to address some of these limitations in medical image segmentation architectures. The proposed approach aims to achieve substantial performance improvements while preserving low computational complexity applicable to both volumetric and 2D contexts.
@@ 3 The Architecture
The structural framework of our proposed U-VixLSTM model is depicted in Fig. 1. It follows the classic $U$-shaped framework, characterized by feature extraction and reconstruction pathways. Although the approach is outlined within the framework of 3D volumetric image processing, it can easily be modified for 2D images by reducing spatial dimension.
The feature extraction arm consists of CNN followed by Vision-xLSTM blocks in the bottleneck, which progressively reduce the size of the feature volumes to generate a high-level abstract representation of the image. The reconstruction path gradually builds the high-dimensional segmentation output, using the low-dimensional contextual representation from the preceding feature extraction path. The output of each convolution block in the feature extraction path is directed to its corresponding counterpart at the same level of the feature reconstruction path, through skip connections. This facilitates integration of feature maps that originate at various levels of abstraction. It ensures a judicious combination of the finer textural details from earlier convolution levels with the coarser semantic information of the deeper levels; thereby, resulting in enhanced context-sensitive predictions.
@@ 3.1 Feature extraction path
This path consists of key components, viz. CNNs for high-level feature learning and Vision-xLSTM for global dependencies.
@@ 3.1.1 High-level features
The volumetric input image $I\in\mathbb{R}^{H\times W\times D\times 1}$ is passed through a series of convolution layers to hierarchically construct an intermediate abstract and high-level representation of the image denoted by $r\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}\times 8C}$. Here $H$, $W$ and $D$ represent the height, width, and depth of the intermediate feature volume, respectively, with $C$ corresponding to the number of channels.
Next, $r$ is divided into $P\times P\times P$ non-overlapping patches. This is followed by the flattening of these patches into 1D vectors to yield a tokenized representation $t\in\mathbb{R}^{N\times(P^{3}\frac{C}{4})}$. Here, $N=(\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4})/P^{3}$ denotes the number of flattened patches with dimension $P^{3}\,\frac{C}{4}$. The flattened patches are then projected into a $Z-$dimensional embedding space, with learnable positional embeddings being added to preserve spatial location information. Mathematically, this is expressed as
$p=[t^{1}\mathbf{K};t^{2}\mathbf{K};\ldots;t^{N}\mathbf{K}]+K_{pos},$ (1)
where $\mathbf{K}\in\mathbb{R}^{(P^{3}\frac{C}{4}\times Z)}$ is the projection matrix and $\mathbf{K_{pos}}\in\mathbb{R}^{N\times Z}$ is the position embedding matrix. $p\in\mathbb{R}^{N\times Z}$ represents the flattened patches.
@@ 3.1.2 Global dependencies
The projected patches are processed by a stack of Vision-xLSTM (ViL) [19] blocks. The structure comprises alternating ViL blocks, with the even-numbered blocks handling patch tokens from the top left to the bottom right, and the odd-numbered ones from the bottom right to the top left. Such bidirectional processing enables ViL blocks to capture robust global dependencies in the input.
The flattened patches, after normalization, are projected onto an embedding space to increase their dimension by a factor of 2. These expanded embeddings are divided into two paths, $x_{mlstm}\in\mathbb{R}^{N\times 2Z}$ and $y\in\mathbb{R}^{N\times 2Z}$ as shown in the detailed representation of the ViL block in Fig. 1. The $x_{mlstm}$ undergoes a 1D causal convolution with SiLU activation [31]. The intermediate result ($\mathbf{X}\in\mathbb{R}^{N\times 2Z}$) is then mapped onto query vector ($q$), key vector ($k$) and value ($v$) vectors, which are similar to the vectors used in the Transformer architecture. These vectors are transmitted to the mLSTM cell. Here, the mLSTM sublayer comprises $n$ parallel attention heads, each equipped with a matrix memory cell. The generation of query, value and keys is mathematically represented as
$\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q}^{T},\quad\mathbf{K}=\mathbf{X}\mathbf{W}_{K}^{T},\quad\mathbf{V}=\mathbf{X}\mathbf{W}_{V}^{T},$ (2)
@@
where $\mathbf{Q}\in\mathbb{R}^{N\times 2Z}$, $\mathbf{K}\in\mathbb{R}^{N\times 2Z}$,$\mathbf{V}\in\mathbb{R}^{N\times 2Z}$ are the query, key and value matrices, respectively. Here, $W_{Q}\in\mathbb{R}^{2Z\times 2Z}$, $W_{K}\in\mathbb{R}^{2Z\times 2Z}$, $W_{V}\in\mathbb{R}^{2Z\times 2Z}$, are learnable weight matrices to generate the query, key, and value vectors.
@@
First the input and forget gate pre-activations, $\tilde{i}\in\mathbb{R}^{N\times 2Z}$ and $\tilde{f}\in\mathbb{R}^{N\times 2Z}$ are calculated by linearly projecting the concatenated $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ matrices as
$\tilde{i}=(\mathbf{W^{I}})^{T}[\mathbf{Q},\mathbf{K},\mathbf{V}]+\mathbf{B},$ (3)
$\tilde{f}=(\mathbf{W^{F}})^{T}[\mathbf{Q},\mathbf{K},\mathbf{V}]+\mathbf{B}.$ (4)
Here, $\mathbf{B}\in\mathbb{R}^{N\times 2Z}$ is the bias matrix. $\mathbf{W^{I}},\mathbf{W^{F}}\in\mathbb{R}^{6Z\times 2Z}$ are the projection matrices for $\tilde{i}$ and $\tilde{f}$ respectively. The pre-activations are used to derive the gate decay matrix $\mathbf{D}\in\mathbb{R}^{N\times 2Z}$ using
$\mathbf{D}=\tilde{i}\oplus l\sigma(\tilde{f}),$ (5)
with $l\sigma$ denoting the log-sigmoid activation function and $\oplus$ represents the element-wise addition operation. The decay matrix $\mathbf{D}$ is further stabilized to ensure that the subsequent exponentiation of $\mathbf{D}$ produces a stable output. The final gate decay matrix output $\hat{\mathbf{D}}\in\mathbb{R}^{N\times 2Z}$ is expressed as
$\hat{\mathbf{D}}=\exp(\mathbf{D}),$ (6)
with $\exp$ representing exponentiation. $\hat{\mathbf{D}}$ regulates the flow of information to the memory matrix. Vectors $\mathbf{Q}$ and $\mathbf{K}$ undergo a dot product multiplication to generate attention scores $\mathbf{S}\in\mathbb{R}^{N\times N}$, analogous to the self-attention mechanism of Transformers. A causal mask $\mathbf{M}\in\mathbb{R}^{N\times N}$ guarantees that attention is directed exclusively from the previous patch to the current patch. This is computed as
$\mathbf{S}=Softmax(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}+\mathbf{M}).$ (7)
The attention score matrix is multiplied by $\hat{\mathbf{D}}$ to get the combination matrix $\mathbf{C}\in\mathbb{R}^{N\times 2Z}$. We have
$\mathbf{C}=\mathbf{S}^{\top}\mathbf{\hat{D}}.$ (8)
Finally, the updated cell state $\tilde{h}_{t}\in\mathbb{R}^{N\times 2Z}$ is computed as
$\tilde{h}_{t}=\mathbf{C}\otimes\mathbf{V}.$ (9)
Here $\otimes$ represents element-wise multiplication operation. The embedding $y$, generated from the second path, is multiplied by $\tilde{h}_{t}$ followed by a down-projection of the final output $o$ onto the $Z$-dimensional embedding space. This is evaluated as
$o=(\mathbf{W^{D}})^{T}(y\otimes\tilde{h}_{t}),$ (10)
with $\mathbf{W^{D}}\in\mathbb{R}^{2Z\times Z}$ being the projection matrix.

### v3

@@ Abstract
The development of efficient segmentation strategies for medical images has evolved from its initial dependence on Convolutional Neural Networks (CNNs) to the current investigation of hybrid models that combine CNNs with Vision Transformers (ViTs). There is an increasing focus on creating architectures that are both high-performing and computationally efficient, capable of being deployed on remote systems with limited resources. Although transformers can capture global dependencies in the input space, they face challenges from the corresponding high computational and storage expenses involved. This research investigates the integration of CNNs with Vision Extended Long Short-Term Memory (Vision-xLSTM)s by introducing the novel U-VixLSTM.
The Vision-xLSTM blocks capture the temporal and global relationships within the patches extracted from the CNN feature maps. The convolutional feature reconstruction path upsamples the output volume from the Vision-xLSTM blocks to produce the segmentation output. Our primary objective is to propose that Vision-xLSTM forms an appropriate backbone for medical image segmentation, offering excellent performance with reduced computational costs. The U-VixLSTM exhibits superior performance compared to the state-of-the-art networks in the publicly available Synapse, ISIC and ACDC datasets. Code provided: https://github.com/duttapallabi2907/U-VixLSTM
@@ 1 Introduction
Artificial intelligence-driven segmentation of medical images helps in the diagnosis, treatment planning, and monitoring of patients in different imaging modalities. There are several challenges in accurately defining target regions from medical image data due to their complex characteristics and variations in anatomical structures. Some common issues include ambiguous boundaries of target organs and pathologies, low contrast, overlapping structures, and inter- and intra-patient variability. Analyzing volumetric medical image data requires a substantial amount of computational resources [1]. These factors lead to the need to develop segmentation algorithms with a high degree of precision and accuracy in the output. Deep learning is widely used in automated medical image segmentation, mainly due to its ability to extract underlying characteristics from input images with minimal human involvement and generalization to unseen samples [2]. CNNs and ViTs are two significant deep learning frameworks used in the design of algorithms for medical image segmentation.
@@
The $U-$Net architecture [3], consisting of symmetric encoder-decoder structures with CNNs, revolutionized the extensive use of CNNs in medical image segmentation. Hierarchical modeling of complex high-level patterns from low-level features and the integration of fine-grained spatial information with coarser abstract features are the factors that are attributed to the success of the $U-$Net. Various well-known architectures have been developed to expand on their accomplishments. These include $U-$Net++ [4], $V-$Net [5], Attention $U-$Net [6], $U-$Net 3+ [7] and LB-UNet [8]. Each introduces distinct modifications to improve performance and address specific difficulties inherent in medical image segmentation.
However, despite the numerous advantages of CNNs, they have inherent limitations. CNNs are incapable of capturing the global context of target anatomical structures with varying sizes due to their limited receptive field. In this context, ViTs became increasingly popular for their ability to capture long-range dependencies between different parts of the input image. This led to exploring combinations of CNNs and Transformers to utilize their generic benefits in creating a robust representation of target structures. ViTs can efficiently learn the overall structural details. In contrast, CNNs focus mainly on local patterns because of their limited field of view [9].
@@
TransUNet [10] pioneered the study of the impact of such a hybrid structure on medical image segmentation. It modeled the global structural dependencies in the intermediate feature volumes, produced from CNN layers, using the self-attention mechanism [11] of ViT blocks. This framework was also adopted into multiple segmentation tasks in the medical domain with the necessary modifications [12], [13], [14] and [15]. UNETR [16], TransAttUNet [17] and UCTransNet [18] are some of the other well-known hybrid CNN-Transformer architectures that integrate CNNs with standard ViT blocks.
The self-attention mechanism, the driving force behind Transformers, suffers from the curse of quadratic complexity. This leads to a huge computational burden and memory requirement. Additionally, the fixed-scale patch processing by ViT blocks across multiple levels lacked multi-scale feature learning capability. This is crucial for efficiently segmenting anatomical structures exhibiting variations in shape and size. Swin Transformers [19] addresses these drawbacks by initially computing self-attention among the patches within non-overlapping windows rather than the entire input. Patch merging layers at every level led to the hierarchical processing of feature maps.
Swin Transformer blocks were leveraged in multiple medical image segmentation models, viz. Swin UNETR [20] and DS-TransUNet [21]. The idea of computing self-attention among patches within a local window was also adopted in [22]. A global token representative of each window was derived and fed to a Gaussian-weighted axial attention module to compute cross-dependencies between these representative tokens. The overall computational complexity got reduced to $\mathcal{O}(n\sqrt{n})$, where $n$ is the total patches. Several other models were developed to reduce the quadratic computational complexity of the self-attention mechanism. The efficient attention mechanism [23] was adopted in [24] to reduce the computational complexity of the self-attention mechanism to $\mathcal{O}(d^{2}n)$, where $d$ is the embedding dimension.
[…]
Mamba-based models lack retaining fine-grained details despite efficient computational and memory complexities while updating the structured memory [29]. This results from the selective structured memory update mechanism that might lead to information loss. Consequently, it leads to inefficient segmentation performance of small and localized anatomical structures. Additionally, Mamba faces difficulties in refining distant past information, which leads to inefficient long-range context correction.
@@
Extended Long-Short Term Memory (xLSTM) [30] has recently emerged as a strong contender to Transformers and Mamba in sequence modeling. It addresses the key limitations of Transformers with its inherent linear computational complexity [$\mathcal{O}(n)$] and constant memory complexity [$\mathcal{O}(1)$] with respect to the length of input sequences. xLSTM extends standard LSTM by incorporating memory structures with higher capacity, in contrast to the scalar memory state of LSTM and compressed state in Mamba. The exponential gating mechanism enhances the ability of the model to revise past information. Vision-xLSTM (ViL) [31] pioneers the application of xLSTMs to computer vision tasks.
@@
Clinical scenarios with limited computing resources require models that can accurately delineate target structures from medical data while constraining computational power and memory requirements. This is essential to guarantee sustainable and cost-effective healthcare solutions. The necessity for computationally efficient and highly accurate segmentation algorithms, coupled with the benefits of xLSTM, motivated us to develop the new U-VixLSTM model for medical image segmentation. Our approach integrates CNNs with ViL for the first time (to our knowledge) for segmentation. The architecture is based on the popular $U-$ shaped framework reported in the literature [32], [33], [34], [35]. Our model is developed in two different versions: one for handling volumetric inputs and the other for handling 2D medical images. Our research contribution is summarized below.
@@
  • CNNs in the feature extraction path initially capture fine-grained textural information and local patterns corresponding to the target anatomical structures from the input image. The ViL block encodes the global context within the intermediate output volumes, as obtained from the CNN layers.
@@
  • The feature reconstruction path upsamples the output from the ViL block to produce the final segmentation output. Skip connections concatenate the feature volume from each level of the feature extraction path to the feature volume of the corresponding level of the feature reconstruction path, to build a robust representation of the target structures.
@@
The remaining sections of the paper are structured in the following manner. Section 2 presents the detailed description of the proposed architecture U-VixLSTM. The experimental results are illustrated in Section 3, along with a comparative study with state-of-the-art networks on publicly accessible datasets to validate the efficacy of U-VixLSTM. Finally, the article is concluded in Section 4.
@@ 2 The Architecture
The structural framework of our proposed U-VixLSTM model is depicted in Fig. 1. It follows the classic $U$-shaped framework, characterized by feature extraction and reconstruction pathways. Although the approach is outlined within the framework of 3D volumetric image processing, it can easily be modified for 2D images by reducing spatial dimensions.
The feature extraction arm has multiple layers of CNNs with ViL blocks in the bottleneck. Each ViL block contains the mLSTM layer [30] to capture long-range dependencies along with temporal awareness. The mLSTM layer employs an exponential gating mechanism to strike a balance between retaining past information and integrating new inputs. The ViL block processes the feature volumes from CNNs to generate an abstract high-level representation of the image. The reconstruction path gradually builds the high-dimensional segmentation output using the contextual representation from the ViL block. The output of each convolution block in the feature extraction path is directed to its corresponding counterpart at the same level of the feature reconstruction path through skip connections. This facilitates the integration of feature maps that originate at various levels of abstraction with the activation maps at the corresponding level of the decoder. It ensures a judicious combination of the finer textural details from earlier convolution levels with the coarser semantic information of the deeper levels, thereby resulting in enhanced context-sensitive predictions.
Figure 1: Architectural framework of U-VixLSTM depicting the input being processed by stacked layers of CNNs and ViL block. The feature representation from the ViL block is upsampled through the feature reconstruction path to obtain the final segmentation output.
@@ 2.1 Feature extraction path
This path consists of key components, viz. CNNs for high-level feature learning and ViL for capturing global dependencies.
@@ 2.1.1 High-level features
The volumetric input image $I\in\mathbb{R}^{H\times W\times D\times 1}$ is passed through a series of convolution layers to hierarchically construct an intermediate abstract and high-level representation of the image denoted by $r\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}\times 8C}$. Here, $H$, $W$ and $D$ represent the height, width, and depth of the intermediate feature volume, respectively, with $C$ corresponding to the number of channels.
Next, $r$ is divided into $P\times P\times P$ non-overlapping patches. This is followed by the flattening of these patches into 1D vectors to yield a tokenized representation $\mathbf{t}\in\mathbb{R}^{N\times(P^{3}\frac{C}{4})}$. Here, $N=(\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4})/P^{3}$ denotes the number of flattened patches with dimension $P^{3}\,\frac{C}{4}$. The flattened patches ($t^{1},t^{2},\ldots,t^{N}$) are then projected into a $Z$-dimensional embedding space, with learnable positional embeddings being added to preserve spatial location information. Mathematically, this is expressed as
$\mathbf{p}=[t^{1}\mathbf{K};t^{2}\mathbf{K};\ldots;t^{N}\mathbf{K}]+\mathbf{K_{pos}},$ (1)
where $\mathbf{K}\in\mathbb{R}^{(P^{3}\frac{C}{4}\times Z)}$ is the projection matrix and $\mathbf{K_{pos}}\in\mathbb{R}^{N\times Z}$ is the position embedding matrix. $\mathbf{p}\in\mathbb{R}^{N\times Z}$ represents the matrix of flattened patches.
Figure 2: Architectural framework of the mLSTM layer to compute inter-patch dependencies.
@@ 2.1.2 Global dependencies
The projected patches are processed by the ViL blocks, with the even-numbered blocks handling patch tokens from the top left to the bottom right, and the odd-numbered ones from the bottom right to the top left. Such bidirectional processing enables ViL to capture robust global dependencies in the input. Inspired by the feed-forward network in [11], the flattened patches, after normalization, are projected onto an embedding space to increase their dimension by a factor of 2. This step enhances the capability of the model to learn complex and non-linear relationships. These expanded embeddings are divided into two paths, $x_{mlstm}\in\mathbb{R}^{N\times 2Z}$ and $y\in\mathbb{R}^{N\times 2Z}$, as shown in Fig. 2. $x_{mlstm}$ is further processed by the mLSTM layer of the ViL block.
The mLSTM layer depicted in Fig. 2 is responsible for modeling the inter-patch dependencies. It is an enhanced variant of LSTM that features a matrix memory cell state $\mathbf{C_{t}}$ rather than a scalar value. At a given time step $t$, $x_{mlstm}$ undergoes a 1D causal convolution with SiLU activation [39]. This step enriches the patch representation by incorporating information from immediate neighbours. The intermediate result ($\mathbf{X_{t}}\in\mathbb{R}^{N\times 2Z}$), which is the current input, is then mapped onto query ($q_{t}$), key ($k_{t}$) and value ($v_{t}$) vectors. The generation of query, value and keys is mathematically represented as
$\mathbf{Q_{t}}=\mathbf{X_{t}}\mathbf{W_{Q_{t}}}^{T},\quad\mathbf{K_{t}}=\mathbf{X_{t}}\mathbf{W_{K_{t}}}^{T},\quad\mathbf{V_{t}}=\mathbf{X_{t}}\mathbf{W_{V_{t}}}^{T},$ (2)
@@
where $\mathbf{Q_{t}}\in\mathbb{R}^{N\times 2Z}$, $\mathbf{K_{t}}\in\mathbb{R}^{N\times 2Z}$,$\mathbf{V_{t}}\in\mathbb{R}^{N\times 2Z}$ are the query, key and value matrices, respectively. Here, $\mathbf{W_{Q_{t}}}\in\mathbb{R}^{2Z\times 2Z}$, $\mathbf{W_{K_{t}}}\in\mathbb{R}^{2Z\times 2Z}$, $\mathbf{W_{V_{t}}}\in\mathbb{R}^{2Z\times 2Z}$ are learnable weight matrices to generate the query, key, and value vectors.
@@
The input and forget gate pre-activations, $\mathbf{i_{t}}\in\mathbb{R}^{N\times 2Z}$ and $\mathbf{f_{t}}\in\mathbb{R}^{N\times 2Z}$ are simultaneously calculated from $\mathbf{X_{t}}$ as
$\mathbf{i_{t}}=\exp((\mathbf{W^{I}})^{T}\mathbf{X_{t}}+\mathbf{B}),$ (3)
$\mathbf{f_{t}}=\exp((\mathbf{W^{F}})^{T}\mathbf{X_{t}}+\mathbf{B}).$ (4)
Here, $\mathbf{B}\in\mathbb{R}^{N\times 2Z}$ is the bias matrix. $\mathbf{W^{I}},\mathbf{W^{F}}\in\mathbb{R}^{2Z\times 2Z}$ are the projection matrices for $\mathbf{i_{t}}$ and $\mathbf{f_{t}}$ respectively. $\exp$ represents the exponentiation operation introduced in xLSTM to overcome the vanishing gradient issue. This enables the model to learn complex relationships among long patch sequences without losing relevant context. The input gate decides which information to store in the current matrix memory cell state $\mathbf{C_{t}}$. The forget gate is responsible for discarding information from the previous cell state at timestamp $t-1$. The cell state $\mathbf{C_{t}}$ is updated as follows:
$\mathbf{C_{t}}=\mathbf{i_{t}}.\mathbf{V_{t}}\mathbf{K_{t}}^{T}+\mathbf{f_{t}}.\mathbf{C_{t-1}}$ (5)
$\mathbf{V_{t}}\mathbf{K_{t}}^{T}$ is the key-value pair representing the matrix of current input information to be stored at time-step $t$. An additional normalizer state $\eta_{t}$ is introduced to tackle the exploding cell state values due to the exponentiation operation. The normalizer state pre-activation is computed as follows:
$\eta_{t}=\mathbf{f_{t}}.\eta_{t-1}+\mathbf{i_{t}}.\mathbf{K_{t}}$ (6)
The normalizer states, until time-step $t-1$, are scaled by $\mathbf{f_{t}}$. This determines the influence of the previous normalizer states on the current normalizer state. The contribution of the current state input $\mathbf{K_{t}}$ is controlled by $\mathbf{i_{t}}$.
Subsequently, the hidden state $h_{t}$ representing the output of the current time-step $t$ is calculated as follows:
$h_{t}=o_{t}[\mathbf{C_{t}/\eta_{t}}]$ (7)
The memory cell state $\mathbf{C_{t}}$ is normalized by $\eta_{t}$ to prevent the exponential growth of cell state values, thereby ensuring numerical stability. $o_{t}=\sigma(\mathbf{W^{o}{{}^{T}}X_{t}}+\mathbf{B})$ represents the output gate activation, which selectively filters the information from normalized $\mathbf{C_{t}}$ to be written to $h_{t}$. The sigmoid activation function, denoted by $\sigma$, squashes $o_{t}$ in the range $[0,1]$. Multiplication ($\odot$) of $h_{t}$ with $y$, as shown in Fig. 2, enhances the patch embedding matrix $y$ with the overall contextual information encoded by $h_{t}$.
ViL processes one patch at a time due to the recurrent structure instead of attending to all the previous patches like ViTs. Thereby, the number of computations remains constant for each patch, which leads to an overall computational complexity of $\mathcal{O}(n)$, where $n$ is the total number of patches. ViL retains a constant-size internal memory throughout, regardless of the total number of patches present, instead of storing the entire set of key-value pairs corresponding to different patches.